OpenStudy (anonymous):

a student bought 1 box of crayonabd 5 reams of paper for $54. she bought 5 boxes of crayon and 3 reams of paper for $50. what is the cost of each box of crayon and each ream of paper?

OpenStudy (anonymous):

let c=cost of crayon and p=cost of paper c+5p=54 5c+3p=50

OpenStudy (anonymous):

how do i solve it

OpenStudy (anonymous):

As a system of equations, by substitution or elimination

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

Multiply the first equation by -5, what do you get

OpenStudy (anonymous):

why do i mulitply it by -5?

OpenStudy (anonymous):

You're going to be adding the two equations together, when you do that you want a variable to be eliminated. c will become -5c and -5c+5c=0

OpenStudy (anonymous):

how did it become negative tho?

OpenStudy (anonymous):

Because you multiply it by -5...

OpenStudy (anonymous):

but dont you have to multipy the 54 by -5 as well?

OpenStudy (anonymous):

Yes you multiply the whole equation by -5. so: -5c-25p=-270 5c+3p=50 now add those equations together

OpenStudy (anonymous):

but where did the -5 come from?

OpenStudy (anonymous):

We want a variable to be eliminated, so you just think of a number that would make that happen.

OpenStudy (anonymous):

so you just make any # up?

OpenStudy (anonymous):

Any number that will make it to where when you add the equations together, one of the variables will cancel out. What do you get when you add them together?